# Controlling coherence between waveguide-coupled quantum dots

### v2

@@ Abstract
Device design.   The device studied in this work is pictured in Fig. 1(e). Two InAs QDs are coupled to a nanobeam waveguide that supports a single guided TE mode. The nanobeam is 280 nm wide and 170 nm thick. The positions of the two QDs are marked by white circles. The distance between the QDs is approximately 20 µm—roughly 70×70$\times$70 × the wavelength of light in the waveguide. The two QDs can be separately addressed using two driving lasers, and emission is collected from the grating outcoupler at the end of the waveguide.
In this work, we use separate Stark tuning of the two sections to control the detuning between the transitions of the QDs. To achieve this, the QDs are located in a split diode structure, allowing different voltages to be applied to different sections of the device. The membrane consists of a p-i-n diode with AlGaAs tunnelling barriers to increase the QD tuning range (see supplementary material [41]). A shallow etch is used to break electrical conductivity through the top p-doped layer of the diode, allowing different voltages to be applied to the separate regions. The positions of this etch are marked in yellow in Fig. 1(e), and the red and blue colouring indicates the two electrically isolated regions. The geometry of the shallow etch through the waveguide is optimised using an inverse design technique to minimise loss; an SEM image of the shallow waveguide etch is inset in Fig. 1(e). We estimate that the electrical isolation etch results in a $<1\%$ transmission loss across a broad bandwidth [41]. At the same time, from numerical simulations we expect this waveguide design to achieve a coupling efficiency ($\beta$ factor) of around 0.8.
@@
Figure 2: Enhancement of emitter decay rate. (a) Comparison of the decay curves of the system with the QDs detuned by (i) 00 µeV, (ii) 1.1 µeV, (iii) 2.2 µeV and (iv) 3.3 µeV. Measurements are performed with (yellow) and without (black) optical gating of QD2. Coloured lines indicate where the intensity drops below the thresholds used in (b) and (c). (b) Detuning dependence of the decay rate with (lighter lines) and without (darker lines) gating of QD2. The results are plotted as the time taken for the emission intensity to drop below three threshold values: $I_{0}/2$ (blue), $I_{0}/e$ (green) and $I_{0}/10$ (red). (c) Theoretical prediction of the results in (b). The shaded regions signify uncertainties in the model due to errors in the $X^{-}_{QD1}$ lifetime estimation (see main text).
@@
The results of the HBT experiments for different detunings can be seen in Fig. 3. For large spectral detunings (top panel) the QDs show uncorrelated emission signified by a drop of the second-order autocorrelation function to 0.5 at zero time-delay. By contrast, when both QDs are tuned into resonance (bottom panel), an anti-dip approaching unity at zero delay time can be observed, which indicates the presence of inter-emitter coherence after the first photon emission, i.e., the preparation of the bright state by the emission process. At intermediate detunings (centre panels), coherence in the emission process can still be observed at detunings far larger than the single-emitter linewidth, and exceeding the region of rate enhancement that has been observed in lifetime measurements (see Fig. 2(a)). The detuning between the emitters then manifests itself in a beating behaviour with a beating frequency corresponding exactly to the QD detuning [41]. Thus, the transition between the correlated emission at zero detuning and uncorrelated emission can be explained by the beating frequency increasing, until it can no longer be resolved by the measurement. Using the measured spectral detuning as input, our theoretical model captures the transition between correlated and independent emission and reproduces the height and frequency of the beating with an estimate of spectral wandering of $\sigma=1.3$ µeV and a pure dephasing rate of $\gamma_{d}=(8.0\pm 1.6)\leavevmode\nobreak\ \textrm{ns}^{-1}$ which are identical across all HBT experiments [41].

### v3

@@ Abstract
Device design. The device studied in this work is pictured in Fig. 1(e). Two InAs QDs are coupled to a nanobeam waveguide that supports a single guided TE mode. The nanobeam is 280 nm wide and 170 nm thick. The positions of the two QDs are marked by white circles. The distance between the QDs is approximately 20 µm—roughly $70\times$ the wavelength of light in the waveguide. The two QDs can be separately addressed using two driving lasers, and emission is collected from the grating outcoupler at the end of the waveguide.
In this work, we use separate Stark tuning of the two sections to control the detuning between the transitions of the QDs. To achieve this, the QDs are located in a split diode structure, allowing different voltages to be applied to different sections of the device. The membrane consists of a p-i-n diode with AlGaAs tunnelling barriers to increase the QD tuning range (see supplementary material [41]). A shallow etch is used to break electrical conductivity through the top p-doped layer of the diode, allowing different voltages to be applied to the separate regions. The positions of this etch are marked in yellow in Fig. 1(e), and the red and blue colouring indicates the two electrically isolated regions. The geometry of the shallow etch through the waveguide is optimised using an inverse design technique to minimise loss; an SEM image of the shallow waveguide etch is inset in Fig. 1(e). We estimate that the electrical isolation etch results in a $1\%$ transmission loss across a broad bandwidth [41]. At the same time, from numerical simulations we expect this waveguide design to achieve a coupling efficiency ($\beta$ factor) of around 0.8.
@@
Figure 2: Enhancement of emitter decay rate. (a) Comparison of the decay curves of the system with the QDs detuned by (i) 0 µeV, (ii) 1.1 µeV, (iii) 2.2 µeV and (iv) 3.3 µeV. Measurements are performed with (yellow) and without (black) optical gating of QD2. Coloured lines indicate where the intensity drops below the thresholds used in (b) and (c). (b) Detuning dependence of the decay rate with (lighter lines) and without (darker lines) gating of QD2. The results are plotted as the time taken for the emission intensity to drop below three threshold values: $I_{0}/2$ (blue), $I_{0}/e$ (green) and $I_{0}/10$ (red). (c) Theoretical prediction of the results in (b). The shaded regions signify uncertainties in the model due to errors in the $X^{-}_{QD1}$ lifetime estimation (see main text).
@@
The results of the HBT experiments for different detunings can be seen in Fig. 3. For large spectral detunings (top panel) the QDs show uncorrelated emission signified by a drop of the second-order autocorrelation function to 0.5 at zero time-delay. By contrast, when both QDs are tuned into resonance (bottom panel), an anti-dip approaching unity at zero delay time can be observed, which indicates the presence of inter-emitter coherence after the first photon emission, i.e., the preparation of the bright state by the emission process. At intermediate detunings (centre panels), coherence in the emission process can still be observed at detunings far larger than the single-emitter linewidth, and exceeding the region of rate enhancement that has been observed in lifetime measurements (see Fig. 2(a)). The detuning between the emitters then manifests itself in a beating behaviour with a beating frequency corresponding exactly to the QD detuning [41]. Thus, the transition between the correlated emission at zero detuning and uncorrelated emission can be explained by the beating frequency increasing, until it can no longer be resolved by the measurement. Using the measured spectral detuning as input, our theoretical model captures the transition between correlated and independent emission and reproduces the height and frequency of the beating with an estimate of spectral wandering of $\sigma=1.3$ µeV and a pure dephasing rate of $\gamma_{d}=(8.0\pm 1.6)\penalty 10000\ \textrm{ns}^{-1}$ which are identical across all HBT experiments [41].